Grade 12Physical Chemistry

At infinite dilution, equivalent conductance for any electrolyte?

Bhagyashri patil
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
 

At infinite dilution, when dissociation is complete, each ion makes a definite contribution towards equivalent conductance of the electrolyte irrespective of the nature of the ion with which it Is associated and the value of equivalent conductance at infinite dilution for any electrolyte is the sum of contribution of its constituent ions", i.e., anions and cations. Thus,

\lambda^\infty _e_q = \lambda^\infty _c +\lambda^\infty _a